OpenStudy (anonymous):

What is 215 to the nearest hundredth?

OpenStudy (anonymous):

200 is the answer

OpenStudy (anonymous):

xlog12=log215 divide both sides by log12 x=log215/log12 so the answer is: x=2.161303737 or rounded to the nearest tenth, it would be: x=2.2

OpenStudy (anonymous):

That would be to the nearest hundred @Bhim

OpenStudy (anonymous):

so its 2.2?

OpenStudy (anonymous):

Heh, @AlexPR787 I don't get it, I think the question is simpler, 215 to the nearest hundredth is 215.00, there are no fractional values for this. @coolaidd Can you clarify the question?

OpenStudy (anonymous):

sorry let me explain

OpenStudy (anonymous):

its 200

OpenStudy (anonymous):

I think there is a misunderstanding: hundredth is http://en.wikipedia.org/wiki/Hundredth

OpenStudy (anonymous):

thats the question..

OpenStudy (anonymous):

it is 200

OpenStudy (anonymous):

Rounding 215 to the nearest hundred would yield 200, but rounding it to the nearest hundredth would mean 215.00, because there are is no fractional part to the value

OpenStudy (anonymous):

so the answer would be?215.00

OpenStudy (anonymous):

Well, I'm wondering if the question may actually mean, what is 215 / 100, to the nearest hundrdeth

OpenStudy (anonymous):

*hundredth

OpenStudy (anonymous):

is that the only thing the question says?

OpenStudy (anonymous):

(1.) If the question is merely, what is 215 to the nearest hundredth, then the answer is 215.00 (2.) If the question is actually seeking, what is 215 divided into 100 parts, 215/100, then what is the value of each part rounded to the nearest hundredth, then you have 215/100 = 2.15, so to the nearest hundredth is 2.15

OpenStudy (anonymous):

just put 2.2

OpenStudy (anonymous):

No, that would be to the nearest tenth

OpenStudy (anonymous):

log_5 (2x-5) =-4 2x - 5 = 5^(-4) = 1/625 2x = 5 1/625 = 3126/625 x = 1563/615

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

well, for that, yeah, looks right

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

do i finish 1563/615?

OpenStudy (anonymous):

If you need to, like if the question requires you to find the answer to the nearest anything

OpenStudy (anonymous):

i got 2.54146341.. what would that be rounded to the hundredth?

OpenStudy (anonymous):

OK, so you want to round to the nearest hundredth, this means you need to consider the value of the thousandth, which is the number in the 3rd decimal place. This requires remember the rules for rounding. If the value in the 3rd decimal place is less than or equal to 5, then you round down, and do nothing to the hundredth value. If the 3rd value is greater than 5, then you round up, meaning you add a one. so if \[2.541\] then round down because of the 1, it's called rounding down because \[2.54 < 2.541\] so if \[2.546\] then round up because of the 1, so \[2.546 becomes 2.54 + .01\] it's called rounding up because \[2.55 > 2.541\]

OpenStudy (anonymous):

so its 2.541?

OpenStudy (anonymous):

to the right of the decimal point you have [tenths, hundredths, thousandths, ten thousandths...] You want to find your answer to the nearest hundredth, so, yes, look at 2.541 The value in the thousandth place is 1, do you round up or down?

OpenStudy (anonymous):

down?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yep, and why ?

OpenStudy (anonymous):

because its below 5?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so 2.541 is right or do i have to do more to it

OpenStudy (anonymous):

Hehe, you have to make to the nearest hundredth, you still have a 1/1000 in there ;)

OpenStudy (anonymous):

so how would it look?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

OK, you want the value with only the hundredth value as the smallest, you want to disregard the other values to the right of the decimal. So, like all the other examples, you would only keep the 2 values to the right of the decimal, so if you have 1.241, you would have 1.24 as the answer if you wanted it to the nearest hundredth....so 2.541 to the nearest hundredth is?

OpenStudy (anonymous):

2.54?

OpenStudy (anonymous):

Yes, exactly

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Sure, I'm going to sign off. I think you can solve most of your questions!